1. Ayla, age 17, is claimed by her parents as a dependent. During 2019, she had interest income from a bank savings account of $2,000 and income from a part-time job of $4,200. Ayla's taxable income is:

a.$4,200 – $4,550 = $0.

b.$6,200 – $4,550 = $1,650.

c.$6,200 – $1,000 = $5,200.

d.$6,200 – $5,700 = $500.

e.None of these choices are correct.

2. In terms of the tax formula applicable to individual taxpayers, which of the following statements, if any, is correct? a.In arriving at taxable income, a taxpayer must choose between the standard deduction and the deduction for qualified business income. b.In arriving at taxable income, a taxpayer must choose between the standard deduction and itemized deductions. c.In arriving at AGI, personal and dependency exemptions are subtracted from gross income. d.The tax formula does not apply if a taxpayer elects to claim the standard deduction. e.None of these choices are correct.
Answer #1

1.

Ayla's taxable income is

Total Income = 4,200 + 2,000 = 6,200
Less: Deduction 4,550

So correct answer is Option b

6,200 - 4,550 = 1,650

2.

The correct answer is

In arriving at taxable income, a taxpayer must choose between the standard deduction and itemized deductions

Option b is correct